# Pure C*-algebras

### v1

@@ 4. Variants of divisibility
This shows that $[a]\in\operatorname{Cu}(A_{\mathcal{U}})$ has (FAD) and (UFAD). Now, using the fact that functional comparison can be restated in terms of comparison in the semigroup ([Rob13, Proposition 2.2.2]), a standard model theoretic argument shows that $[a]\in\operatorname{Cu}(A)$ also has (FAD) and (UFAD); see [FHL+21, Proposition 8.1.6] and also [APRT24, Theorem 10.5]. ∎
@@ 5. Pure C*-algebras
In this section, we show that every $(m,n)$-pure $C^{*}$-algebra is pure; see Theorem 5.8.
@@
Given $m,n\in{\mathbb{N}}$, we say that a $\mathrm{Cu}$-semigroup $S$ is $(m,n)$-pure if $S$ has $m$-comparison (Definition 3.1) and is $n$-almost divisible (Definition 4.1). We say that $S$ is pure if it is $(0,0)$-pure.
A $\mathrm{Cu}$-semigroup $S$ has strict comparison if, whenever $x,y\in S$ with $x\leq\infty y$ satisfy that $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ normalized at $y$, one has $x\leq y$.
A $C^{*}$-algebra is $(m,n)$-pure if its Cuntz semigroup is. Similarly, a $C^{*}$-algebra is pure if it is $(0,0)$-pure. We also say that a $C^{*}$-algebra has strict comparison of positive elements (by $2$-quasitraces) provided its Cuntz semigroup has strict comparison.
It was proved in [ERS11, Proposition 6.2] that $S$ has strict comparison if and only if it is almost unperforated.
@@
Every $\mathcal{Z}$-stable $C^{*}$-algebra is pure. Every pure $C^{*}$-algebra has strict comparison of positive elements.
@@
Let $A$ be a $\mathcal{Z}$-stable $C^{*}$-algebra. It was shown in [Rør04, Theorem 4.5] that $\operatorname{Cu}(A)$ is almost unperforated, and thus by the comment above $A$ has strict comparison of positive elements. The argument used to prove almost unperforation can be used to show almost divisibility. For a proof, see e.g. [APT11, Theorem 5.35]. ∎
Remark 5.3.
It is perhaps worth pointing out that the results in [Rør04] deal with the uncomplete Cuntz semigroup $\mathrm{W}(A)$. This makes no difference regarding almost unperforation, since $\operatorname{Cu}(A)$ is a completion of $\mathrm{W}(A)$ (see [APT18, Theorem 3.2.8]), hence $\mathrm{W}(A)$ is almost unperforated if and only if $\operatorname{Cu}(A)$ is.
@@
Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup satisfying (O5) and (O7). Assume that $S$ has locally bounded comparison amplitude, and also that is upper and lower functionally almost divisible. Then $S$ is pure.
@@
Therefore, let $x^{\prime}\ll x$ and pick $x^{\prime\prime}\in S$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$. Since we have $\widehat{x}\leq\tfrac{n}{n+1}\widehat{y}$, we may choose $\gamma,\gamma^{\prime}$ such that $\tfrac{n}{n+1}<\gamma<\gamma^{\prime}<1$. Using Proposition 2.3 (i) one gets $\widehat{x^{\prime\prime}}\ll\gamma\widehat{y}$, which allows us to pick $y^{\prime}\in S$ such that
@@
Take $y^{\prime\prime}$ such that $y^{\prime}\ll y^{\prime\prime}\ll y$, and choose $d\in{\mathbb{N}}$ such that $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. Since $(S,\Sigma)$ has locally bounded comparison amplitude, we obtain $N=N(\gamma^{\prime},d)\in{\mathbb{N}}$ such that
$\widehat{v}\leq\gamma^{\prime}\widehat{w}\quad\Rightarrow\quad v\leq_{N}w,\quad\text{ for all }v,w\in\Sigma^{(d)}.$
Let $M\in{\mathbb{N}}$ be such that $M\geq 2$, $M\geq N$ and $\gamma\tfrac{M+1}{M-1}\leq\gamma^{\prime}$. Using (UFAD) for $x^{\prime}\ll x^{\prime\prime}$ and $M-1$, we obtain $v\in S$ satisfying
$v\leq x^{\prime\prime},\quad(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}},\,\,\,\text{ and }\,\,\,x^{\prime}\leq Mv.$
Applying (LFAD) for $y^{\prime}\ll y^{\prime\prime}$ and $M$, one gets $w\in S$ such that
$Mw\leq y^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{y^{\prime}}\leq(M+1)\widehat{w}.$
@@
$(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}}\leq\gamma\widehat{y^{\prime}}\leq\gamma(M+1)\widehat{w},$
which implies, by our choice of $M$, that $\widehat{v}\leq\gamma^{\prime}\widehat{w}$.
Note that both $v$ and $w$ belong to $\Sigma^{(d)}$, since $v\leq x^{\prime\prime},Mw\leq y^{\prime\prime}$, and $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. We deduce that $v\leq_{N}w$ and, since $M\geq N$, we get $Mv\leq Mw$.
Hence,
$x^{\prime}\leq Mv\leq Mw\leq y^{\prime\prime}\leq y,$
@@
To show that $S$ is almost divisible, let $k\geq 2$ and take $x^{\prime},x\in S$ such that $x^{\prime}\ll x$. Since $S$ satisfies (LFAD), we may apply Lemma 4.4(ii) to obtain $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
It follows that $x^{\prime}\leq\infty y$; see, for example, [APRT22, Lemma 6.6 (iii)]. We also have that $\lambda(x^{\prime})<\lambda((k+1)y)$ for every functional normalized at $y$. Since $S$ is almost unperforated, we obtain $x^{\prime}\leq(k+1)y$ by [ERS11, Proposition 6.2]. ∎
@@
Let $A$ be a $C^{*}$-algebra that has locally bounded comparison amplitude and is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $A$ is pure.
@@
Since $\operatorname{Cu}(A)$ is $n$-almost divisible, Theorem 4.7 gives that $\operatorname{Cu}(A)$ has (LFAD) and (UFAD), and the result follows from Proposition 5.6. ∎
@@
If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.

### v3

@@ 4. Variants of divisibility
A finite collection of Cuntz subequivalences in $A_{\mathcal{U}}$ can be ‘lifted’ to $A$ simultaneously up to a cut-down (see, for example, the proof of [APRT24, Theorem 10.5]). Thus, using that $x^{\prime}\ll x=[a]$, we find $y,z\in\operatorname{Cu}(A)$ such that
$ky\leq x,\quad sNx^{\prime}\leq s(Nky+x),$
and
$z\leq x,\quad t(N-1)kz\leq tNx,\quad x^{\prime}\leq kz$
in $\operatorname{Cu}(A)$. Passing to functionals, we can remove the multiplication by $s$ and $t$, and we obtain $N\widehat{x^{\prime}}\leq Nk\widehat{y}+\widehat{x}$ and $(N-1)k\widehat{z}\leq N\widehat{x}$, which shows that $y$ and $z$ have the desired properties. ∎
It was proved in [APRT22, Theorem 3.8] that, if $A$ is a separable $C^{*}$-algebra of stable rank one, then $\operatorname{Cu}(A)$ is an inf-semilattice ordered semigroup. That is, for each $x,y\in\operatorname{Cu}(A)$ their greatest lower bound $x\wedge y$ exists in $\operatorname{Cu}(A)$ and addition is distributive over the meet operation:
$(x\wedge y)+z=(x+z)\wedge(y+z)$
for all $x,y,z\in\operatorname{Cu}(A)$. This will be used below.
Proposition 4.10.
Every nowhere scattered $C^{*}$-algebra with stable rank one is functionally divisible.
Proof.
Let us first assume that every separable, nowhere scattered $C^{*}$-algebra with stable rank one is functionally divisible, and let $A$ be an arbitrary nowhere scattered $C^{*}$-algebra with stable rank one. To show that $A$ is functionally divisible, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$ and let $\varepsilon\in(0,1)$ be given. We need to find $y,z\in\operatorname{Cu}(A)$ such that
$ky\leq x,\quad(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y},\quad z\leq x,\quad(1-\varepsilon)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
Using that stable rank one and nowhere scatteredness satisfy the Löwenheim-Skolem condition ([TV24a, Proposition 4.11]) and applying [TV21, Proposition 6.1], we can find a separable sub-$C^{*}$-algebra $B\subseteq A$ such that $B$ is nowhere scattered, has stable rank one, and such that the inclusion $B\to A$ induces an order-embedding $\operatorname{Cu}(B)\to\operatorname{Cu}(A)$ whose image contains $x^{\prime}$ and $x$. Viewing $x^{\prime}$ and $x$ as elements in $\operatorname{Cu}(B)$, and using that $B$ is functionally divisible, we find $y$ and $z$ with the desired properties in $\operatorname{Cu}(B)$, and thus in $\operatorname{Cu}(A)$. (See also [APRT22, Lemma 9.2].)
By the argument above, we may assume that $A$ is a separable, nowhere scattered $C^{*}$-algebra $A$ with stable rank one, and we need to show that $A$ is functionally divisible. We first verify condition (iii) in Proposition 4.5, which then implies that $\operatorname{Cu}(A)$ satisfies condition (L).
Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, and let $\varepsilon>0$. We need to find $y\in\operatorname{Cu}(A)$ such that
$2y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y}.$
Using the solution to the rank problem (see Theorem 2.4) we can find $u\in\operatorname{Cu}(A)$ such that $\widehat{u}=\tfrac{1}{2}\widehat{x}$. Set $v:=u\wedge x$, which exists by the comments preceding this proposition. Using [APRT22, Theorem 6.12] at the second step, we have
$\widehat{v}=\widehat{u\wedge x}=\widehat{u}\wedge\widehat{x}=\frac{1}{2}\widehat{x}.$
Using that $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible by the results in [APRT22, Section 5] (see also Theorem 7.1 and Proposition 7.3 in [TV23]), we can apply [AVTV23, Theorem 5.10] to find a soft element $w\in\operatorname{Cu}(A)$ such that
$w\leq v,\,\,\,\text{ and }\,\,\,\widehat{w}=\widehat{v}.$
By Proposition 2.3(i), we have
$(1-\varepsilon)\frac{1}{2}\widehat{x^{\prime}}\ll\frac{1}{2}\widehat{x}=\widehat{w},$
which allows us to choose $w^{\prime\prime},w^{\prime}\in\operatorname{Cu}(A)$ such that
$w^{\prime\prime}\ll w^{\prime}\ll w,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\frac{1}{2}\widehat{x^{\prime}}\leq\widehat{w^{\prime\prime}}.$
Since $w$ is soft, there exists by [TV24b, Proposition 4.6] some $t\in\operatorname{Cu}(A)$ such that
$w^{\prime}+t\ll w,\,\,\,\text{ and }\,\,\,w^{\prime}\ll\infty t.$
In particular, $w^{\prime}+t\leq v\leq x$.
Choose $t^{\prime}\in\operatorname{Cu}(A)$ such that
$t^{\prime}\ll t,\,\,\,\text{ and }\,\,\,w^{\prime}\ll\infty t^{\prime}.$
Applying (O5) for $w^{\prime}+t\leq x$ and $w^{\prime\prime}\ll w^{\prime}$ and $t^{\prime}\ll t$, we obtain $c\in\operatorname{Cu}(A)$ such that
$w^{\prime\prime}+c\leq x\leq w^{\prime}+c,\,\,\,\text{ and }\,\,\,t^{\prime}\ll c.$
Let us show that $\tfrac{1}{2}\widehat{x}\leq\widehat{c}$. Since $w^{\prime}\ll\infty t^{\prime}$, there is $m\geq 1$ such that $w^{\prime}\leq mt^{\prime}$, and thus $x\leq w^{\prime}+c\leq mt^{\prime}+t^{\prime}=(m+1)t^{\prime}$. Therefore, if $\lambda\in F(\operatorname{Cu}(A))$ satisfies $\lambda(x)=\infty$, then $\lambda(c)=\infty\geq\tfrac{1}{2}\lambda(x)$. On the other hand, if $\lambda\in F(\operatorname{Cu}(A))$ satisfies $\lambda(x)<\infty$, then using that $w^{\prime}\leq w$ and $\widehat{w}=\tfrac{1}{2}\widehat{x}$, we get
$\lambda(c)\geq\lambda(x)-\lambda(w^{\prime})\geq\lambda(x)-\frac{1}{2}\lambda(x)=\frac{1}{2}\lambda(x).$
Set $y:=w^{\prime\prime}\wedge c$. Using that $(1-\varepsilon)\tfrac{1}{2}\widehat{x^{\prime}}\leq\widehat{w^{\prime\prime}}$ and $\tfrac{1}{2}\widehat{x}\leq\widehat{c}$, and using [APRT22, Theorem 6.12] at the second step, we have
$(1-\varepsilon)\tfrac{1}{2}\widehat{x^{\prime}}\leq\widehat{w^{\prime\prime}}\wedge\widehat{c}=\widehat{w^{\prime\prime}\wedge c}=\widehat{y},$
and thus
$2y\leq w^{\prime\prime}+c\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y},$
as desired.
Next, to verify condition (U) from Definition 4.3, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon\in(0,1)$. We need to find $z\in S$ such that
$z\leq x,\quad(1-\varepsilon)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
Set $s:=(1-\varepsilon)^{-1}\tfrac{1}{k}$. If $s\geq 1$ (that is, if $(1-\varepsilon)k\leq 1$), then we can use $z:=x$. Thus, we may assume that $s<1$. For $f\in L(F(\operatorname{Cu}(A)))$, let $\alpha(f)\in\operatorname{Cu}(A)$ be as in Theorem 2.4, hence $\widehat{\alpha(f)}=f$. Set
$z:=x\wedge\alpha\left(s\widehat{x}\right).$
Then $z\leq x$. Further, using that $s<1$, and using [APRT22, Theorem 6.12], we have
$\widehat{z}=\widehat{x}\wedge s\widehat{x}=s\widehat{x}=\frac{(1-\varepsilon)^{-1}}{k}\widehat{x}$
and thus $(1-\varepsilon)k\widehat{z}=\widehat{x}$.
Finally, let us verify that $x^{\prime}\leq kz$. In fact we will see that $x\leq kz$. Since $\operatorname{Cu}(A)$ is inf-semilattice ordered, we have
$kz=k\big{(}x\wedge\alpha(s\widehat{x})\big{)}=\bigwedge_{i=0}^{k}\big{(}ix+(k-i)\alpha(s\widehat{x})\big{)}.$
Clearly $x\leq ix+(k-i)\alpha(s\widehat{x})$ when $i>0$. For $i=0$, first note that $ks-1>0$ and therefore $\widehat{x}\leq\infty(ks-1)\widehat{x}$, which allows us to apply the partial additivity of $\alpha$ stated in Theorem 2.4. Hence, we get
$x\leq x+\alpha\big{(}(ks-1)\widehat{x}\big{)}=\alpha\big{(}\widehat{x}+(ks-1)\widehat{x}\big{)}=\alpha(ks\widehat{x})=k\alpha(s\widehat{x}),$
and thus $x\leq kz$, as desired. ∎
Proposition 4.11.
Let $A$ be a simple, unital, non-elementary $C^{*}$-algebra with a unique normalized quasitrace. Then $\operatorname{Cu}(A)$ is functionally divisible.
Proof.
By assumption, there exists a unique functional $\lambda\colon\operatorname{Cu}(A)\to[0,\infty]$ satisfying $\lambda([1])=1$. Note that any two elements $x,y\in\operatorname{Cu}(A)$ satisfy $\widehat{x}\leq\widehat{y}$ if and only if $\lambda(x)\leq\lambda(y)$.
We first verify condition (iii) in Proposition 4.5, which then implies that $\operatorname{Cu}(A)$ satisfies condition (L) in Definition 4.3. Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, and let $\varepsilon\in(0,1)$. We need to find $y\in\operatorname{Cu}(A)$ such that
$2y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\lambda(x^{\prime})\leq 2\lambda(y).$
If $x^{\prime}=0$, then we can use $y=0$. Therefore, we may from now on assume that $x^{\prime}\neq 0$. Then $\lambda(x^{\prime})\in(0,\infty)$. Since $A$ is non-elementary, we can pick $v\in\operatorname{Cu}(A)$ such that $\lambda(v)=\tfrac{1}{2}\lambda(x^{\prime})$. We have
$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\frac{1}{2}\lambda(x^{\prime})\leq\lambda(x),\lambda(v).$
Using that $\operatorname{Cu}(A)$ satisfies Edwards’ condition for $\lambda$ (see Theorem 4.7 and Remark 4.2(3) in [Thi20], and [APRT21]), we obtain $w\in\operatorname{Cu}(A)$ such that
$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(w),\,\,\,\text{ and }\,\,\,w\leq v,x.$
Choose $w^{\prime\prime},w^{\prime}\in\operatorname{Cu}(A)$ such that
$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(w^{\prime\prime}),\,\,\,\text{ and }\,\,\,w^{\prime\prime}\ll w^{\prime}\ll w.$
Applying (O5) for $w^{\prime\prime}\ll w^{\prime}\leq x$, we obtain $c\in\operatorname{Cu}(A)$ such that
$w^{\prime\prime}+c\leq x\leq w^{\prime}+c.$
Using that $\lambda(x^{\prime})\leq\lambda(x)$ and $\lambda(w^{\prime})\leq\lambda(w)\leq\lambda(v)=\tfrac{1}{2}\lambda(x^{\prime})$, we get
$\frac{1}{2}\lambda(x^{\prime})=\lambda(x^{\prime})-\frac{1}{2}\lambda(x^{\prime})\leq\lambda(x)-\lambda(w^{\prime})\leq\lambda(c).$
Then
$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(w^{\prime\prime}),\,\,\,\text{ and }\,\,\,(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\frac{1}{2}\lambda(x^{\prime})\leq\lambda(c).$
Applying Edwards’ condition again, we obtain $y\in\operatorname{Cu}(A)$ such that
$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(y),\,\,\,\text{ and }\,\,\,y\leq w^{\prime\prime},c.$
Then $2y\leq w^{\prime\prime}+c\leq x$. Thus $y$ has the desired properties.
Next, to verify condition (U) of Definition 4.3, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon\in(0,1)$. We need to find $z\in\operatorname{Cu}(A)$ such that
$z\leq x,\quad(1-\varepsilon)k\lambda(z)\leq\lambda(x),\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
Choose $\delta>0$ such that
$1+(2k-1)\delta\leq(1-\varepsilon)^{-1}.$
Then pick $v^{\prime\prime},v^{\prime},v\in\operatorname{Cu}(A)$ such that
$x^{\prime}\ll v^{\prime\prime}\ll v^{\prime}\ll v\ll x,\,\,\,\text{ and }\,\,\,\lambda(v)\leq(1+\delta)\lambda(v^{\prime\prime}).$
We have already verified condition (L) of Definition 4.3. Applied for $v^{\prime}\ll v$, for $k$ and for $\tfrac{\delta}{2}$, we obtain $y\in\operatorname{Cu}(A)$ such that
$ky\leq v,\,\,\,\text{ and }\,\,\,\left(1-\frac{\delta}{2}\right)\lambda(v^{\prime})\leq k\lambda(y).$
By Proposition 2.3(i), we have $(1-\delta)\widehat{v^{\prime\prime}}\ll(1-\tfrac{\delta}{2})\widehat{v^{\prime}}$ in $\operatorname{Lsc}(F(\operatorname{Cu}(A)))$, which allows us to choose $y^{\prime}\in\operatorname{Cu}(A)$ such that
$y^{\prime}\ll y,\,\,\,\text{ and }\,\,\,(1-\delta)\lambda(v^{\prime\prime})\leq k\lambda(y^{\prime}).$
Applying Lemma 2.1 for $ky\leq v$ and $y^{\prime}\ll y$, we obtain $z\in\operatorname{Cu}(A)$ such that
$(k-1)y^{\prime}+z\leq v\leq(k-1)y+z,\,\,\,\text{ and }\,\,\,v\leq kz.$
Using that $\lambda(y^{\prime})$ is finite, we get
$k\lambda(z)$ $\leq k\lambda(v)-(k-1)k\lambda(y^{\prime})$
$\leq k(1+\delta)\lambda(v^{\prime\prime})-(k-1)(1-\delta)\lambda(v^{\prime\prime})$
$=(1+(2k-1)\delta)\lambda(v^{\prime\prime})\leq(1-\varepsilon)^{-1}\lambda(v^{\prime\prime}),$
and thus
$(1-\varepsilon)k\lambda(z)\leq\lambda(v^{\prime\prime})\leq\lambda(x).$
We further have
$z\leq v\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq v\leq kz,$
which shows that $z$ has the claimed properties. ∎
@@ 5. Pure C*-algebras
This section is devoted to proving Theorems A and C. We actually show that a $C^{*}$-algebra is pure if, and only if, it has controlled comparison and is functionally divisible; if, and only if, it is $(m,n)$-pure for some $m$ and $n$. Building on the results from Section 4 concerning automatic functional divisibility, we show that a $C^{*}$-algebra with controlled comparison is pure if it is either nowhere scattered with real rank zero or stable rank one, or also if it is simple, unital, non-elementary with a unique normalized quasitrace.
@@
Given $m,n\in{\mathbb{N}}$, we say that a $\mathrm{Cu}$-semigroup is $(m,n)$-pure if it has $m$-comparison and is $n$-almost divisible. We say that a $\mathrm{Cu}$-semigroup is pure if it is $(0,0)$-pure.
A $C^{*}$-algebra is $(m,n)$-pure if its Cuntz semigroup is. Similarly, a $C^{*}$-algebra is pure if it is $(0,0)$-pure.
@@
Every $\mathcal{Z}$-stable $C^{*}$-algebra is pure. Every pure $C^{*}$-algebra has the Global Glimm Property and strict comparison (of positive elements by quasitraces).
@@
Let $A$ be a $\mathcal{Z}$-stable $C^{*}$-algebra. Then the (classical, uncomplete) Cuntz semigroup $W(A)$ is almost unperforated by [Rør04, Theorem 4.5]. Further, by [APT11, Theorem 5.35], for every $x\in W(A)$ and every $k\geq 1$ there exists $y\in W(A)$ such that $ky\leq x\leq(k+1)y$. Using that $W(A)$ is order-dense in $\operatorname{Cu}(A)$ (see [APT18, Theorem 3.2.8]), it follows that $\operatorname{Cu}(A)$ is almost unperforated and almost divisible (see also Paragraph 4.2), and thus pure.
By definition, every pure $C^{*}$-algebra is almost divisible, and therefore $(2,\omega)$-divisible, which by [TV23, Theorem 3.6] is equivalent to the Global Glimm Property. Further, every pure $C^{*}$-algebra has 00-comparison, and thus enjoys strict comparison as noted in Paragraph 3.4. ∎
@@
Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup satisfying (O5). Assume that $(S,\Sigma)$ has controlled comparison and is functionally divisible. Then $S$ is pure.
@@
Therefore, let $x^{\prime}\in S$ with $x^{\prime}\ll x$, and pick $x^{\prime\prime}\in S$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$. Since we have $\widehat{x}\leq\tfrac{n}{n+1}\widehat{y}$, we may choose $\gamma,\gamma^{\prime}$ such that $\tfrac{n}{n+1}<\gamma<\gamma^{\prime}<1$. Using Proposition 2.3(i) one gets $\widehat{x^{\prime\prime}}\ll\gamma\widehat{y}$, which allows us to pick $y^{\prime}\in S$ such that
@@
Take $y^{\prime\prime}$ such that $y^{\prime}\ll y^{\prime\prime}\ll y$, and choose $d\in{\mathbb{N}}$ such that $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. Since $(S,\Sigma)$ has controlled comparison, we obtain $N=N(\gamma^{\prime},d)\in{\mathbb{N}}$ such that
$\widehat{v}\leq\gamma^{\prime}\widehat{w}\quad\text{ implies }\quad v\leq_{N}w,\quad\text{ for all }v,w\in\Sigma^{(d)}.$
Pick $\varepsilon>0$ such that $\gamma\leq(1-\varepsilon)^{2}\gamma^{\prime}$. Applying condition (U) in Definition 4.3 for $x^{\prime}\ll x^{\prime\prime}$ and $N$, we obtain $v\in S$ satisfying
$v\leq x^{\prime\prime},\quad(1-\varepsilon)N\widehat{v}\leq\widehat{x^{\prime\prime}},\,\,\,\text{ and }\,\,\,x^{\prime}\leq Nv.$
Similarly, applying condition (L) in Definition 4.3 for $y^{\prime}\ll y^{\prime\prime}$ and $N$, one gets $w\in S$ such that
$Nw\leq y^{\prime\prime},\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{y^{\prime}}\leq N\widehat{w}.$
@@
$(1-\varepsilon)^{2}N\widehat{v}\leq(1-\varepsilon)\widehat{x^{\prime\prime}}\leq(1-\varepsilon)\gamma\widehat{y^{\prime}}\leq\gamma N\widehat{w}\leq(1-\varepsilon)^{2}N\gamma^{\prime}\widehat{w},$
which implies that $\widehat{v}\leq\gamma^{\prime}\widehat{w}$.
Note that both $v$ and $w$ belong to $\Sigma^{(d)}$, since $v\leq x^{\prime\prime},Nw\leq y^{\prime\prime}$, and $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. It follows that $v\leq_{N}w$, and thus
$x^{\prime}\leq Nv\leq Nw\leq y^{\prime\prime}\leq y,$
@@
To show that $S$ is almost divisible, let $x^{\prime}\ll x$ in $S$, and let $k\in{\mathbb{N}}$. Pick $\varepsilon\in(0,1)$ such that $(1-\varepsilon)^{-1}k<k+1$. Applying condition (L) in Definition 4.3 for $x^{\prime\prime}\ll x$, and for $k$ and $\varepsilon$, we obtain $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}.$
It follows that $x^{\prime}\leq\infty y$; see, for example, [APRT22, Lemma 6.6(iii)]. Since $(1-\varepsilon)^{-1}k<k+1$, we also have that $\lambda(x^{\prime})<\lambda((k+1)y)$ for every functional normalized at $y$. Since $S$ is almost unperforated, we obtain $x^{\prime}\leq(k+1)y$ by [ERS11, Proposition 6.2]. ∎
Remark 5.6.
It is possible to adapt the proof of Proposition 5.5 so that instead of (U) it suffices to assume the condition ($U_{2^{m}}$) from Remark 4.6 for all $m$. Namely, the following stronger form of Proposition 5.5 is available: If a scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ satisfies (O5) and has controlled comparison, and for every $x^{\prime}\ll x$ in $S$ and every $\varepsilon\in(0,1)$ there exist $y,z\in S$ such that
$2y\leq x,\quad(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y},\quad z\leq x,\quad(1-\varepsilon)2\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq 2z,$
then $S$ is pure.
@@
Given a $C^{*}$-algebra $A$, the following statements are equivalent:
  1. (i)
    $A$ is pure, that is, $A$ has strict comparison and is almost divisible.
  2. (ii)
    $A$ is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$.
  3. (iii)
    $A$ has controlled comparison and is functionally divisible.
@@
It is clear that (i) implies (ii). Further, (ii) implies (iii) since $m$-comparison implies controlled comparison by Proposition 3.7, and $n$-almost divisibility implies functional divisibility by Proposition 4.9. Finally, (iii) implies (i) by Proposition 5.5 since the Cuntz semigroup of every $C^{*}$-algebra satisfies (O5). ∎
Remark 5.8.
If follows from Theorem 5.7 (see also Paragraph 4.2) that $A$ is $(m,n)$-pure if, and only if, $\operatorname{Cu}(A)$ has $m$-comparison and is $n$-almost divisible in the sense of [Win12, Definition 3.5], that is, given $x\in\operatorname{Cu}(A)$ and $k\geq 1$ , there is $y\in\operatorname{Cu}(A)$ such that $ky\leq x\leq(n+1)(k+1)y$.
Remark 5.9.
Theorem 5.7 shows that the combination of weak forms of comparison and divisibility implies strong forms of comparison and divisibility. One may wonder if this already holds individually for comparison and divisibility.
For the comparison properties, this is not the case. Indeed, there exist $C^{*}$-algebras that have controlled comparison (or even $m$-comparison for some $m\geq 1$) that do not have 00-comparison. For example, if $X=[0,1]^{5}$ is the five-dimensional cube, then $C(X)$ has nuclear dimension at most five, and therefore has $5$-comparison by [Rob11, Theorem 1.3], but it does not have 00-comparison since the radius of comparison is at least one by [EN13, Theorem 1.1].
Similarly, we expect that the answer to the following question is positive.
Question 5.10.
Does there exist a $C^{*}$-algebra that is functionally divisible (or even $n$-almost divisible for some $n$) but not almost divisible?
Proposition 5.11.
Let $A$ be a functionally divisible $C^{*}$-algebra. Then the following statements are equivalent:
  1. (i)
    $A$ is pure.
  2. (ii)
    $A$ has strict comparison (of positive elements by quasitraces).
  3. (iii)
    $A$ has $m$-comparison for some $m$.
  4. (iv)
    $A$ has controlled comparison.
Proof.
By definition, pureness implies 00-comparison, which is equivalent to strict comparison by Paragraph 3.4. This shows that (i) implies (ii), which in turn implies (iii). Further, (iii) implies (iv) by Proposition 3.7. Finally, (iv) implies (i) by Theorem 5.7. ∎
Combining Proposition 5.11 with the results proving functional divisibility from Section 4, we can now verify pureness in various settings:
@@
Let $A$ be a nowhere scattered $C^{*}$-algebra that has real rank zero or stable rank one. Assume that $A$ has controlled comparison (for example, $m$-comparison for some $m$). Then $A$ is pure.